OpenStudy (anonymous):

The area of a regular octagon is 35 cm2. What is the area of a regular octagon with sides five times as long? A. 625 cm2 B. 875 cm2 C. 175 cm2 D. 245 cm2

OpenStudy (anonymous):

I thought it was C but I got it wrong =(

OpenStudy (anonymous):

It really should be C... Weird..

OpenStudy (anonymous):

Maybe it is an answer key issue- it is an autoscored exam. Do you think it is worth it to dispute it? Or could we be wrong?

OpenStudy (anonymous):

Hold on... Let me work it again.

OpenStudy (fifciol):

I got B

OpenStudy (anonymous):

Yes, it is B because you need to remember that 5 needs to be 5x5 for the entire figure. So it would be 35x25.

OpenStudy (anonymous):

Thank you guys so much! Sometimes the simplest part is the easiest to forget. It makes perfect sense now. THANK YOU!

OpenStudy (fifciol):

\[\frac{ 8a^2\sqrt3 }{4 }=35\] \[\frac{ 8(5a)^2\sqrt3 }{ 4 }=x\]

OpenStudy (anonymous):

No problem. Glad I remember something. :P I'm weak in math.

OpenStudy (anonymous):

A parallelogram has sides measuring 23.8 m and 35.3 m. The height corresponding to the 23.8-m base is 9.9 m. Find the height, to the nearest tenth of a meter, corresponding to the 35.3-m base. A. 8317.4 m B. 14.7 m C. 6.7 m D. none of these

OpenStudy (anonymous):

This is the last one I am having trouble with. I thought it was B but got it wrong.

OpenStudy (anonymous):

@lissa_bug I am the same- math kicks my butt! =(

OpenStudy (fifciol):

the area of parallelogram is : side times height corresponding to that side, so 23,8*9,9=35,3*h

OpenStudy (fifciol):

what is h then?

OpenStudy (anonymous):

h=6.67

OpenStudy (anonymous):

so it would be C

OpenStudy (fifciol):

yes

OpenStudy (anonymous):

Thank you again! Not sure how I messed that one up. But you explain things in easy to understand terms.

OpenStudy (fifciol):

yw:)